OpenStudy (anonymous):

What is the vertex for the graph of y = 2x2 + 12x + 10? (−3, −8) (−3, −44) (3, 64) (3, 58)

OpenStudy (dumbcow):

do you want the short answer or long answer :)

OpenStudy (anonymous):

Short would be good for now please!:)

OpenStudy (anonymous):

(-3,-8)

OpenStudy (dumbcow):

ok formula for "x" component of vertex is: x = -b/2a (where a,b,c are coefficients of equation) x = -12/4 = -3 plug in x =-3 to get "y" component

OpenStudy (anonymous):

just differentiate the above eq.. at the point of vertex the slope is 0 for a quadratic function equate ur answer to 0 to get x component and substitute it in ur function to get the y component

OpenStudy (dumbcow):

haha i dont think calculus has been learned yet

OpenStudy (anonymous):

the formula u gave is itself derived from calculus

OpenStudy (dumbcow):

i know that but students learn formula first then how its derived later....thats how it always works

OpenStudy (aravindg):

he formula can also be got by completing the square

OpenStudy (anonymous):

ok

OpenStudy (dumbcow):

finding vertex of parabola is an algebra or pre-calc course

OpenStudy (aravindg):

*The